### v1

# The Gaussian-Multinoulli Restricted Boltzmann Machine A Potts Model Extension of the GRBM

Abstract
Many real-world tasks, from associative memory to symbolic reasoning, demand discrete, structured representations that standard continuous latent models struggle to express naturally. We introduce the Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM), a generative energy-based model that extends the Gaussian-Bernoulli RBM (GB-RBM) by replacing binary hidden units with $q$-state Potts variables. This modification enables a combinatorially richer latent space and supports learning over multivalued, interpretable latent concepts. We formally derive GM-RBM’s energy function, learning dynamics, and conditional distributions, showing that it preserves tractable inference and training through contrastive divergence. Empirically, we demonstrate that GM-RBMs model complex multimodal distributions more effectively than binary RBMs, outperforming them on tasks involving analogical recall and structured memory. Our results highlight GM-RBMs as a scalable framework for discrete latent inference with enhanced expressiveness and interoperability.
1 Introduction
Restricted Boltzmann Machines (RBMs) constitute a class of undirected energy-based models defined on a bipartite graph between visible units $\mathbf{v}$ and latent units $\mathbf{h}$(hidden). Their tractable block Gibbs updates [6] and scalable approximate inference via contrastive divergence enable efficient training on high-dimensional datasets [1, 6]. In the binary hidden configuration, an RBM reduces to an Ising model where hidden units $h_{j}\in{-1,+1}$ interact pairwise through couplings $W_{ij}$, yielding an energy landscape analogous to classical spin systems [14]. The Gaussian–Bernoulli RBM (GB-RBM) augments this formulation by replacing these visible variables with Gaussian potentials, while preserving binary hidden spins [7]. Although GB-RBMs capture continuous inputs such as image intensities or sensor readings, the binary latent constraints limit their expressivity when representing multimodal or discrete feature correlations [4, 15] .
The $q$-state Potts model generalizes the Ising model by allowing each spin $s_{j}$ to occupy one of $q$ discrete states, enriching the model’s capacity to encode complex interaction patterns and phase transitions [22, 32]. Potts-type energy functions have demonstrated efficacy in image segmentation, combinatorial optimization, community detection, and categorical embedding tasks, where capturing higher-order discrete structures is key [12]. Building on these principles, we introduce the Gaussian–Multinoulli Restricted Boltzmann Machine (GM-RBM), which integrates Gaussian-distributed visible units with a hidden layer of $q$-state Potts spins. This GB-RBM-like architecture provides a clear path to benchmarking and showcasing the power of Potts-based energy functions.
The Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM) we propose has a small change in the underlying representation of hidden nodes that has a profound improvement on model performance. While prior work has addressed mixed-variate inputs using Gaussian-Bernoulli structures [26], convolutional and tiled weight-sharing schemes for spatial hierarchies [10], and implicit mixture models of RBMs for multimodal density estimation [20], these approaches retain binary or sigmoidal-based latent representations, limiting their ability to capture categorical variation at the feature level. In contrast, the GM-RBM replaces binary hidden units with $q$-state Potts (Multinoulli) variables, enabling each hidden unit to express one of several discrete latent modes rather than mere presence or absence. This yields a latent space that scales as $q^{m}$ (with $m$ hidden units), and facilitates more structured, interpretable, and compositional generative behavior. Unlike energy-based models that treat discrete latents as auxiliary sampling artifacts [4], or continuous latent embeddings for mixed data [25], the GM-RBM directly leverages categorical latent dynamics to build rich distributions over continuous-valued visible units. This hybrid formulation opens new avenues for modeling structured analog data, such as trajectories, natural images, or biophysical time series, with discrete control over generative features.
2 Background and Model Overview
2.1 Background and Motivation
Restricted Boltzmann Machines (RBMs) have played a foundational role in energy-based modeling and unsupervised representation learning. In particular, Gaussian-Bernoulli RBMs (GB-RBMs), which pair real-valued Gaussian visible units with binary latent variables, have been adopted for modeling continuous input data such as images and signals [15]. However, while GB-RBMs are capable of learning low-level features, they face inherent limitations in their ability to capture structured, discrete latent representations.
The root of this limitation lies in the mismatch between the binary latent space and the Gaussian visible manifold. The Gaussian prior enforces a continuous energy landscape in the input space, which, when combined with binary latent units, leads to the learning of clustered, overlapping encodings. These encodings tend to coalesce into fuzzy representations that lack the expressiveness needed for capturing discrete abstractions, such as object parts, symbolic states, or multimodal feature variants.
Moreover, the binary latent units in GB-RBMs act as on/off switches, providing limited control over the diversity and structure of the hidden space. This bottleneck hampers the model’s ability to disentangle categorical or compositional structure, particularly when the data exhibits rich semantic variation.
2.2 Motivating Discrete Latent Structure: Potts Units
In contrast, Potts models, which extend binary variables to q-state categorical variables, offer a compelling mechanism for reintroducing structured discrete latent representations into energy-based models. By allowing each hidden unit to express one of multiple mutually exclusive states, Potts units can represent discrete modes, templates, or symbolic features directly. This increases the combinatorial capacity of the model from $2^{m}$ (binary) to $q^{m}$ (categorical), enabling richer and more interpretable feature representations.
Importantly, Potts units also induce a sharper factorization of latent structure, allowing for the emergence of compositional encodings where different latent units capture orthogonal or complementary concepts. This categorical structure aligns more naturally with real-world semantics and improves both generalization and controllability.
2.3 The Gaussian-Multinoulli RBM (GM-RBM)
To address the limitations of GB-RBMs and exploit the representational advantages of Potts variables, we introduce the Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM). In this architecture, the visible layer consists of Gaussian-valued units, modeling real-valued data such as sensor signals or pixel intensities while the hidden layer consists of Multinoulli (Potts) units, each of which selects one of q discrete latent states.
Each Potts state corresponds to a different “template” or filter that contributes to the energy of the model, allowing a single hidden unit to encode latent modes or feature classes rather than just presence/absence. The GM-RBM thus combines the smooth representational power of Gaussian inputs with the symbolic richness of discrete latent categories.
This hybrid construction restores the model’s ability to capture structured, discrete latent features while retaining flexibility in modeling continuous-valued observations. As we show, the GM-RBM can learn more interpretable latent spaces, support combinatorial recombination of features, and exhibit sharper mode separation in the visible manifold than its Gaussian-Bernoulli counterpart.
3 Theoretical Foundations of the Gaussian-Multinoulli RBM (GM-RBM)
The Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM) extends the classical Gaussian-Bernoulli RBM by replacing binary latent units with discrete $q$-state Potts (Multinoulli) variables. This modification introduces a combinatorially structured latent space capable of expressing discrete modes while modeling continuous-valued data in the visible layer. In this section, we develop a mathematical understanding of how this architecture supports richer and more structured representations.
3.1 Energy Function and Architecture
The GM-RBM defines a joint energy over visible variables $v\in\mathbb{R}^{n}$ and hidden Potts variables $h=(h_{1},\dots,h_{m})$, where each $h_{j}\in\{1,\dots,q\}$ selects a discrete state for hidden unit $j$:
$E(v,h)=\frac{1}{2}\sum_{i=1}^{n}(v_{i}-b_{i})^{2}-\sum_{j=1}^{m}c_{j,h_{j}}-\sum_{i=1}^{n}\sum_{j=1}^{m}W_{ij}^{(h_{j})}v_{i}$ (1)
This can be re-expressed in terms of the conditional Gaussian mean $\mu(h)$ as:
$E(v,h)=\frac{1}{2}\|v-\mu(h)\|^{2}+\text{const}(h),\quad\text{where}\quad\mu(h)=b+\sum_{j=1}^{m}W_{j}^{(h_{j})}$ (2)
Each hidden unit contributes a state-specific template vector $W_{j}^{(h_{j})}$, and the visible layer is modeled as a Gaussian centered at the compositional mean $\mu(h)$:
$P(v\mid h)=\mathcal{N}(\mu(h),I)$ (3)
3.2 Compositionality and Factored Representation
The mean vector $\mu(h)$ is a sum over learned templates selected by the Potts states:
$\mu(h)=b+\sum_{j=1}^{m}W_{j}^{(h_{j})}$ (4)
This creates a structured and factored representation in visible space. Each hidden unit encodes one of $q$ possible modes, and the overall visible configuration is a compositional sum of these latent components. This modularity enables reuse of latent features across samples, disentangled latent representation, and combinatorial expressivity across $q^{m}$ configurations
The set of all compositional means $\mu(h)$ forms a structured lattice in $\mathbb{R}^{n}$. The resulting energy landscape has the following properties:
  • Locally linear: Each mode defines a Gaussian with fixed covariance
  • Globally discrete: The mode structure is defined by a combinatorial Potts lattice
  • Modular: Latent units contribute independently to the visible pattern
This structure allows GM-RBMs to model multimodal distributions over continuous space with precise symbolic control via latent discrete selections.
4 Hetero-associative Memory
Hetero-associative memory refers to a system’s capability to learn paired associations between distinct patterns, allowing the retrieval of a target pattern (e.g., a response word) when presented with a corresponding stimulus (e.g., a cue word) [11, 19]. This concept, rooted in cognitive modeling and neural computation, was further explored by Hinton in 1981 [2] and later employed in language-related tasks using Gaussian-Bernoulli Restricted Boltzmann Machines (GB-RBMs) [27]. However, the binary hidden units of GB-RBMs impose a hard ceiling on representational capacity.
Across the board, GM-RBMs deliver consistently higher recall than GB-RBMs, with gains amplifying at larger $q$ and corpus sizes even when the number of parameters were identical between the two models. It is important to note that this performance gain is obtained although the GM-RBM was trained using only Gibbs sampling, while the GB-RBM used a stronger and more expensive hybrid Gibbs-Langevin Sampling.
4.1 Experimental Setup
4.1.1 Dataset
We replicated the experimental setup in Tsutsui and Hagiwara, constructing a word-pair dataset representing conceptual relationships (e.g., "apple is-a fruit") [27]. We randomly selected pairs from WordNet [28], excluding compound or incomplete entries and sampled 500-3,000 word pairs to create smaller datasets for training and testing scalability.
Each word pair is treated as a directional association task and encoded as a concatenated embedding vector [17], mirroring the hetero-associative memory objective of recalling a target concept from a given stimulus.
4.1.2 Preprocessing
We trained a 200-dimensional Continuous Bag of Words (CBOW) Word2Vec model [18] on the word-pair dataset (100 iterations, window size 5, no frequency cutoff). We normalized each vector dimension to zero mean and unit variance to mitigate small-magnitude issues. For each stimulus-response pair, embeddings were concatenated to form a 400-dimensional input vector for the RBM’s visible layer.
4.1.3 Computational Setup
We used a compute node (Intel Xeon Gold 6154 ×2, 512 GB RAM, NVIDIA Tesla P40) running Red Hat Enterprise Linux 8. Models were trained with CUDA-accelerated PyTorch (v1.13) on a single GPU. Experiments were automated via a modular framework for data loading, configuration, visualization, and checkpointing.
4.1.4 Training
Our GM-RBM learned to associate stimulus–response word embeddings. Datasets comprised semantically related word pairs (e.g., doctor–nurse, sun–light). Following a similar procedure to Tsutsui and Hagiwara [27], we trained a CBOW Word2Vec model [23] on these two-word sentences to capture in-domain semantics, producing 200-dimensional embeddings (100 iterations, window size 5, no frequency cutoff).
We normalized each embedding dimension to zero mean and unit variance to avoid numerical instability. The stimulus and response embeddings were concatenated into a 400-dimensional visible layer vector.
We trained the GB-RBM using contrastive divergence with a two-step Gibbs burn-in, employing the more expensive Gibbs–Langevin sampling procedure. In contrast, the GM-RBM variant relied solely on standard Gibbs sampling, reducing computational overhead. Notably, in the special case of $q=2$, the GM-RBM formulation becomes almost equivalent to the GB-RBM: the only difference is that in the GB-RBM, the two weight matrices are essentially negatives of one another. Hidden-unit counts were scaled inversely with the number of Potts states to keep capacity constant. Training used Adam (LR = $10^{-4}$) with mini-batches of 64. We evaluated recall by inferring responses via Gibbs sampling and selecting the nearest neighbor; accuracy was the percentage of correct matches.
We stopped training early when recall accuracy reached 0.98 on the validation set, when the standard deviation of validation accuracy over 20 checkpoints fell below 0.01, or when no improvement was observed for 10 consecutive checkpoints.
4.2 Results
To isolate the effect of the Potts hidden units, we did two key tests. The first was a parameter-matched comparison where the total number of weights was kept equal while $q$ was increased. This ensured that any observed gains in recall accuracy arise purely from the richer Potts representation. The second was a similar sweep where $q$ was held constant while the number of hidden units themselves was increased. Both sweeps were done across varying dataset sizes, measuring the accuracy of the recall in pairs embedded in Word2Vec [18].
Each marker in Figures 1 and 2 denotes an independently trained model using identical early-stopping criteria. Error bars have been omitted as the observed performance gains were uniformly large, making statistical uncertainty negligible for the comparative trends shown.
For data shown in Figure 1, it is important to note that the total number of parameters was held constant i.e. hidden layers were decreased proportionally to the cardinality of the Potts state $q$ (See 1).
It is also important to note $q=2$ for the GM-RBM is a case of the Potts nodes where weights do not have to be constrained i.e. negatives of one another. The original GB-RBM outperforms the GM-RBM for $q=2$, possibly due to the lack of this constraint. Despite the GM-RBMs being trained using the simpler Gibbs update, the $q=4,6,8,10$ GM-RBM models all vastly outperform the GB-RBM, which used the more expensive Gibbs-Langevin update [15].
Since constraining the total number of parameters to be identical might obfuscate the potential of both models, we swept the number of hidden nodes and dataset sizes as shown in Figure 2. It can be clearly seen that GM-RBM $q=2$ and the GB-RBM both fail when $N>2000$, while the GM-RBM $q=4$ maintains its retrieval accuracy.
4.2.1 Parameter matched q sweep
Figure 1: Retrieval accuracy versus the number of associative pairs for different numbers of Potts states (q) in a parameter-matched GB-RBM and GM-RBM setup
To empirically showcase the direct effect of the additional Potts states, we held the total number of model parameters constant while varying the number of Potts states $q$, isolating the effect of the state space’s structure on hetero-associative performance. Figure 1 plots retrieval accuracy against the number of associative pairs in the training set for each $q$. For the binary case $q=2$, the GB-RBM with Gibbs–Langevin sampling maintains near-perfect accuracy at small dataset sizes but collapses sharply beyond 1000 pairs, whereas the GM-RBM using only Gibbs sampling degrades more rapidly. In contrast, models with higher state cardinality ($q=4,6,8,10$) sustain almost perfect retrieval up to roughly 1200–1500 pairs and exhibit a more gradual decline as a function of the dataset size($N$). Notably, $q=10$ consistently outperforms lower-$q$ configurations even for large $N$.
To keep the model capacity fixed across different Potts-state configurations, we solve for the number of hidden units $n_{h}$ by dividing the total budget of weight parameters $n_{w}$ by the size of each hidden Potts spin $q$. Since each hidden unit with $q$ Potts states contributes $q$ weight vectors of length equal to the number of visible units $n_{v}$, then number of hidden units is given by
$n_{h}\texttt{=}\lceil\frac{n_{w}}{n_{v}}\times q\ \rceil$ (5)
We chose $\mathrm{n_{w}}=800{,}000$, ensuring that the binary case $q=2$ matched the optimal hidden-unit count of 1,000 reported by Tsutsui and Hagiwara [27]for the GB-RBM. This ensures that, as we increase the number of Potts states $k$, we reduce the number of hidden units proportionally so that every model has the same total number of parameters.
Table 1: Potts-state $q$ versus number of hidden units
$\,q=2$ $\,q=4$ $\,q=6$ $\,q=8$ $\,q=10$
Hidden Units 1000 500 333 250 200
This result demonstrates increasing the number of discrete states enhances memory robustness under constant parameter constraints. By allocating representational capacity across more Potts states, the model tolerates larger associative loads, suggesting an optimal trade-off point around moderate state sizes (e.g., $q=6$ or $8$) for typical dataset scales in the 2000–2500 pair range.
4.2.2 Hidden nodes sweep
Figure 2: Retrieval accuracy on a semantic hetero-associative memory task as a function of hidden-layer size for different model variants and dataset sizes. (a) GB-RBM (Gibbs–Langevin Update), (b) GM-RBM with $q=2$ (Gibbs Update), and (c) GM-RBM with $q=4$ (Gibbs Update). Each curve corresponds to a different number of training word-pair examples ($N$).
In order to showcase how varying the hidden-layer dimensionality affects retrieval performance across dataset sizes, we swept the number of hidden units from 500 to 3000 while fixing the state cardinality $q$. Figure 2 presents retrieval accuracy curves for three models: a standard Gaussian–Bernoulli RBM (GB-RBM), a GM-RBM with $q=2$, and a GM-RBM with $q=4$). Each curve represents a different dataset size ($N=500,1000,1500,2000,2500,3000$ associative word pairs).
For the binary GM-RBM ($q=2$), accuracy is near-perfect at small loads but degrades sharply as $N$ increases, recovering only when hidden units exceed 1500. In contrast, the Potts-based GM-RBM with $q=4$ maintains over 90% accuracy across all dataset sizes with just 1000 hidden units. The GB-RBM baseline requires roughly 2500 hidden units to achieve similar performance at large scales (e.g., $N=2500$).
These findings demonstrate a clear trade-off between hidden-layer dimensionality and state complexity: increasing $q$ substantially lowers the hidden-unit requirement for robust associative recall, yielding a more parameter-efficient memory architecture.
5 Auto-associative memory
As a "proof-of-concept" to demonstrate the generative capability of our GM-RBM, we performed a replication of the generative experiments from the original GB-RBM paper [15]. On MNIST, we trained for 500 epochs, and on CelebA for 100 epochs. Starting from i.i.d. Gaussian noise in the visible layer and running 1,000 steps of Gibbs sampling, we obtained samples (see Figure 3) that, by visual inspection, showcasing the GM-RBM’s ability to still produce high-quality generative outputs.
5.1 Experimental setup
5.1.1 Datasets
We target two key datasets to showcase the GM-RBM’s generative abilities: MNIST (28×28 grayscale handwritten digits) [13] and CelebA (Center-cropped and resized RGB facial images at 64×64 resolution)  [16].
5.1.2 Preprocessing
All image datasets are normalized to zero mean and unit variance per channel [15]. CelebA images are center-cropped to 140×140 before downsampling to 64×64 [16]. The GMM datasets are constructed using predefined priors and covariance structures, allowing for exact density visualization and sampling [15]. For each dataset, we generate training and evaluation splits using standard protocols.
5.1.3 Computational resources
Experiments were performed on a dedicated compute node equipped with dual Intel Xeon Gold 5218 CPUs, 512 GB of RAM, and eight NVIDIA RTX 6000 GPUs connected via NVLink. All models were trained using CUDA-accelerated PyTorch (v1.13+) with single-GPU execution unless otherwise specified.
5.2 Results
(a) Selected samples from MNIST
(b) Generation of CelebA images from random noise
Figure 3: Sampled results from GM-RBM of MNIST and CelebA datatsets
We observe that the $q=4$ GM-RBM begins to generate visually identifiable face/digit samples with an order of magnitude lower number of epochs compared to the GB-RBM trained with Gibbs–Langevin sampling (See Table 2. Although a fully controlled head-to-head comparison under identical hyperparameter budgets (see Table 2) requires further investigation (See Section 6.1, it is important to emphasize two key distinctions. Our GM-RBM relies solely on conventional Gibbs sampling, whereas Hinton’s GB-RBM employs the more computationally expensive Gibbs–Langevin procedure [15]. This ability is most likely due to what the rapid mixing of Potts models compared to binary Boltzmann machines [22, 31, 32]. In addition, we maintain the same total number of parameters while increasing the number of Potts states.
This reduction in both sampling complexity and overall training budget underscores the richer latent representational capacity afforded by Potts-style hidden units, which can capture more nuanced multimodal structure with fewer resources. The stark difference in hyperparameters was chosen specifically to highlight the GM-RBM’s ability to achieve high-quality samples under a dramatically reduced resource budget, showcasing its practical efficiency.
Table 2: Hyperparameter settings for GB-RBM and GM-RBM on MNIST and CelebA
Hyperparameter MNIST CelebA
GB-RBM  [15] GM-RBM GB-RBM [15] GM-RBM
Number of states 2 4 2 4
Sampling style Gibbs–Langevin Gibbs Gibbs–Langevin Gibbs
Hidden nodes 4096 2048 10000 5000
Visible nodes 784 784 3072 3072
Epochs trained 3000 500 10000 100
6 Limitations and future work
While the Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM) demonstrates improved expressiveness and recall performance over the binary GB-RBM, key limitations and future work remain.
6.1 Limitations of the current model
1. Sampling constraints: In the current implementation, only Gibbs sampling is supported for the Potts-based GM-RBM. While this allows for tractable inference over discrete states, the absence of Langevin or hybrid Gibbs-Langevin samplers limits potential gains in sample quality and mixing speed. As $q$ increases, the discrete state space grows, and relying solely on categorical sampling may lead to slower convergence or poor exploration of the energy landscape in more complex settings.
2. Evaluation domain: The current evaluation focuses on word pair associations using Word2Vec embeddings. The generalization of GM-RBMs to other domains, such as vision, audio, or time series data, remains an open question and will require further validation. These limitations present valuable directions for future research, including scaling GM-RBMs to deeper architectures, optimizing discrete sampling methods, and exploring broader applications beyond associative memory tasks.
3. Epochs and training: Currently, the GM-RBM is trained for a limited number of epochs (500 for MNIST, 100 for CelebA) to achieve initial convergence. While this rapid training demonstrates efficiency, it may not capture long-term stability or the potential benefits of extended training schedules. Optimal epoch budgets, early-stopping criteria, and the impact of prolonged training on sample diversity and model robustness are areas for further exploration.
6.2 Broader impact
6.2.1 Potts units in energy transformers
Energy Transformers (ETs) minimize a global energy via recurrent updates and have shown success in tasks such as image inpainting [9]. Replacing their binary hidden units with $q$-state Potts variables raises the latent capacity from $2^{H}$ to $q^{H}$, yielding:
$E_{\mathrm{mem}}(h)=-\sum_{\mu=1}^{M}\sum_{i=1}^{H}\delta(h_{i},\xi_{i}^{\mu})$ (6)
where ${\xi^{\mu}}$ are stored patterns. Larger $q$ reduces overlap between attractors, improving convergence selectivity. Embedding Gaussian–Potts RBMs into ETs may boost robustness and facilitate neuromorphic implementations; an empirical study of these hybrids is ongoing.
6.2.2 Extending DBMs to continuous data
Deep Boltzmann Machines (DBMs) stack Bernoulli–Bernoulli RBMs to learn hierarchical features via layer-wise Gibbs sampling and contrastive divergence [3, 24]. Their reliance on binary visibles limits real-valued inputs. As a front end, we propose a Gaussian–Potts energy:
$E(v,h)=\sum_{i=1}^{D}\frac{(v_{i}-a_{i})^{2}}{2\sigma_{i}^{2}}-\sum_{i,j}W_{ij}(v_{i},h_{j})-\sum_{j}b_{h_{j}},$ (7)
with $h\_j\in{1,\dots,q}$. This encoder maps $\mathbb{R}^{D}$ into a $q^{H}$-state latent space, preserving DBM training while improving stability and expressiveness.
6.3 Bernoulli-Variable models and Potts benefits
Many foundational generative architectures rely on binary or Bernoulli units, including classical models like Bernoulli RBMs [6], Gaussian–Bernoulli RBMs [15], Deep Belief Networks [24], Deep Boltzmann Machines [3, 24], and Hopfield Networks. More recent advancements leveraging binary sampling mechanisms include Variational Autoencoders (VAEs) [5] with Bernoulli latents, Gumbel–Softmax VAEs [21] that provide differentiable relaxations for discrete sampling, Autoregressive Models such as PixelCNN [30] and PixelRNN [29] for sequential binary pixel generation, Binarized Neural Networks (BNNs) that constrain weights and activations to binary values, Latent Bernoulli Autoencoders [5] designed explicitly for compact binary representations, Spiking Neural Networks (SNNs) employing discrete spike events for communication, and Discrete Denoising Diffusion Models (D3PMs) [8] that extend diffusion processes to discrete data through multinomial transition kernels. Upgrading these binary units to Potts variables further increases representational granularity and sharpens attractor basins, significantly reducing interference among stored patterns and enabling richer discrete structures, ultimately enhancing the capability and expressiveness of associative and generative models.
6.4 Hardware-Efficient Implementation
Many of these models can be deployed with minimal overhead by leveraging sparse or event-driven representations (e.g., SPAD arrays [31]) and one-hot encoding of discrete states. Binary and Potts variables map naturally to bitwise logic and lookup tables, enabling FPGA, ASIC, and neuromorphic implementations with low memory footprint and energy consumption.

### v2

@@
-The Gaussian-Multinoulli Restricted Boltzmann Machine A Potts Model Extension of the GRBM
+The Gaussian-Multinoulli Restricted Boltzmann Machine: A Potts Model Extension of the GRBM
 Abstract
-Many real-world tasks, from associative memory to symbolic reasoning, demand discrete, structured representations that standard continuous latent models struggle to express naturally.
-We introduce the Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM), a generative energy-based model that extends the Gaussian-Bernoulli RBM (GB-RBM) by replacing binary hidden units with $q$-state Potts variables.
-This modification enables a combinatorially richer latent space and supports learning over multivalued, interpretable latent concepts.
-We formally derive GM-RBM’s energy function, learning dynamics, and conditional distributions, showing that it preserves tractable inference and training through contrastive divergence.
-Empirically, we demonstrate that GM-RBMs model complex multimodal distributions more effectively than binary RBMs, outperforming them on tasks involving analogical recall and structured memory.
-Our results highlight GM-RBMs as a scalable framework for discrete latent inference with enhanced expressiveness and interoperability.
+Many real-world tasks, from associative memory to symbolic reasoning, benefit from discrete, structured representations that standard continuous latent models can struggle to express.
+We introduce the Gaussian–Multinoulli Restricted Boltzmann Machine (GM-RBM), a generative energy-based model that extends the Gaussian–Bernoulli RBM (GB-RBM) by replacing binary hidden units with $q$-state categorical (Potts) units, yielding a richer latent state space for multivalued concepts.
+We provide a self-contained derivation of the energy, conditional distributions, and learning rules, and detail practical training choices (contrastive divergence with temperature annealing and intra-slot diversity constraints) that avoid state collapse.
+To separate architectural effects from sheer latent capacity, we evaluate under both capacity-matched and parameter-matched setups, comparing GM-RBM with GB-RBM configured to have the same number of possible latent assignments.
+On analogical recall and structured memory benchmarks, GM-RBM achieves competitive, and in several regimes, improved recall at equal capacity with comparable training cost, despite using only Gibbs updates.
+The discrete $q$-ary formulation is also amenable to efficient implementation.
+These results clarify when categorical hidden units provide a simple, scalable alternative to binary latents for discrete inference within tractable RBMs.
 1 Introduction
-Restricted Boltzmann Machines (RBMs) constitute a class of undirected energy-based models defined on a bipartite graph between visible units $\mathbf{v}$ and latent units $\mathbf{h}$(hidden).
-Their tractable block Gibbs updates [6] and scalable approximate inference via contrastive divergence enable efficient training on high-dimensional datasets [1, 6].
-In the binary hidden configuration, an RBM reduces to an Ising model where hidden units $h_{j}\in{-1,+1}$ interact pairwise through couplings $W_{ij}$, yielding an energy landscape analogous to classical spin systems [14].
-The Gaussian–Bernoulli RBM (GB-RBM) augments this formulation by replacing these visible variables with Gaussian potentials, while preserving binary hidden spins [7].
-Although GB-RBMs capture continuous inputs such as image intensities or sensor readings, the binary latent constraints limit their expressivity when representing multimodal or discrete feature correlations [4, 15] .
-The $q$-state Potts model generalizes the Ising model by allowing each spin $s_{j}$ to occupy one of $q$ discrete states, enriching the model’s capacity to encode complex interaction patterns and phase transitions [22, 32].
-Potts-type energy functions have demonstrated efficacy in image segmentation, combinatorial optimization, community detection, and categorical embedding tasks, where capturing higher-order discrete structures is key [12].
-Building on these principles, we introduce the Gaussian–Multinoulli Restricted Boltzmann Machine (GM-RBM), which integrates Gaussian-distributed visible units with a hidden layer of $q$-state Potts spins.
-This GB-RBM-like architecture provides a clear path to benchmarking and showcasing the power of Potts-based energy functions.
-The Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM) we propose has a small change in the underlying representation of hidden nodes that has a profound improvement on model performance.
-While prior work has addressed mixed-variate inputs using Gaussian-Bernoulli structures [26], convolutional and tiled weight-sharing schemes for spatial hierarchies [10], and implicit mixture models of RBMs for multimodal density estimation [20], these approaches retain binary or sigmoidal-based latent representations, limiting their ability to capture categorical variation at the feature level.
-In contrast, the GM-RBM replaces binary hidden units with $q$-state Potts (Multinoulli) variables, enabling each hidden unit to express one of several discrete latent modes rather than mere presence or absence.
-This yields a latent space that scales as $q^{m}$ (with $m$ hidden units), and facilitates more structured, interpretable, and compositional generative behavior.
-Unlike energy-based models that treat discrete latents as auxiliary sampling artifacts [4], or continuous latent embeddings for mixed data [25], the GM-RBM directly leverages categorical latent dynamics to build rich distributions over continuous-valued visible units.
-This hybrid formulation opens new avenues for modeling structured analog data, such as trajectories, natural images, or biophysical time series, with discrete control over generative features.
-2 Background and Model Overview
-2.1 Background and Motivation
-Restricted Boltzmann Machines (RBMs) have played a foundational role in energy-based modeling and unsupervised representation learning.
-In particular, Gaussian-Bernoulli RBMs (GB-RBMs), which pair real-valued Gaussian visible units with binary latent variables, have been adopted for modeling continuous input data such as images and signals [15].
-However, while GB-RBMs are capable of learning low-level features, they face inherent limitations in their ability to capture structured, discrete latent representations.
-The root of this limitation lies in the mismatch between the binary latent space and the Gaussian visible manifold.
-The Gaussian prior enforces a continuous energy landscape in the input space, which, when combined with binary latent units, leads to the learning of clustered, overlapping encodings.
-These encodings tend to coalesce into fuzzy representations that lack the expressiveness needed for capturing discrete abstractions, such as object parts, symbolic states, or multimodal feature variants.
-Moreover, the binary latent units in GB-RBMs act as on/off switches, providing limited control over the diversity and structure of the hidden space.
-This bottleneck hampers the model’s ability to disentangle categorical or compositional structure, particularly when the data exhibits rich semantic variation.
-2.2 Motivating Discrete Latent Structure: Potts Units
-In contrast, Potts models, which extend binary variables to q-state categorical variables, offer a compelling mechanism for reintroducing structured discrete latent representations into energy-based models.
-By allowing each hidden unit to express one of multiple mutually exclusive states, Potts units can represent discrete modes, templates, or symbolic features directly.
-This increases the combinatorial capacity of the model from $2^{m}$ (binary) to $q^{m}$ (categorical), enabling richer and more interpretable feature representations.
-Importantly, Potts units also induce a sharper factorization of latent structure, allowing for the emergence of compositional encodings where different latent units capture orthogonal or complementary concepts.
-This categorical structure aligns more naturally with real-world semantics and improves both generalization and controllability.
-2.3 The Gaussian-Multinoulli RBM (GM-RBM)
-To address the limitations of GB-RBMs and exploit the representational advantages of Potts variables, we introduce the Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM).
-In this architecture, the visible layer consists of Gaussian-valued units, modeling real-valued data such as sensor signals or pixel intensities while the hidden layer consists of Multinoulli (Potts) units, each of which selects one of q discrete latent states.
-Each Potts state corresponds to a different “template” or filter that contributes to the energy of the model, allowing a single hidden unit to encode latent modes or feature classes rather than just presence/absence.
-The GM-RBM thus combines the smooth representational power of Gaussian inputs with the symbolic richness of discrete latent categories.
-This hybrid construction restores the model’s ability to capture structured, discrete latent features while retaining flexibility in modeling continuous-valued observations.
-As we show, the GM-RBM can learn more interpretable latent spaces, support combinatorial recombination of features, and exhibit sharper mode separation in the visible manifold than its Gaussian-Bernoulli counterpart.
-3 Theoretical Foundations of the Gaussian-Multinoulli RBM (GM-RBM)
-The Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM) extends the classical Gaussian-Bernoulli RBM by replacing binary latent units with discrete $q$-state Potts (Multinoulli) variables.
-This modification introduces a combinatorially structured latent space capable of expressing discrete modes while modeling continuous-valued data in the visible layer.
-In this section, we develop a mathematical understanding of how this architecture supports richer and more structured representations.
-3.1 Energy Function and Architecture
-The GM-RBM defines a joint energy over visible variables $v\in\mathbb{R}^{n}$ and hidden Potts variables $h=(h_{1},\dots,h_{m})$, where each $h_{j}\in\{1,\dots,q\}$ selects a discrete state for hidden unit $j$:
-$E(v,h)=\frac{1}{2}\sum_{i=1}^{n}(v_{i}-b_{i})^{2}-\sum_{j=1}^{m}c_{j,h_{j}}-\sum_{i=1}^{n}\sum_{j=1}^{m}W_{ij}^{(h_{j})}v_{i}$ (1)
-This can be re-expressed in terms of the conditional Gaussian mean $\mu(h)$ as:
-$E(v,h)=\frac{1}{2}\|v-\mu(h)\|^{2}+\text{const}(h),\quad\text{where}\quad\mu(h)=b+\sum_{j=1}^{m}W_{j}^{(h_{j})}$ (2)
-Each hidden unit contributes a state-specific template vector $W_{j}^{(h_{j})}$, and the visible layer is modeled as a Gaussian centered at the compositional mean $\mu(h)$:
-$P(v\mid h)=\mathcal{N}(\mu(h),I)$ (3)
-3.2 Compositionality and Factored Representation
-The mean vector $\mu(h)$ is a sum over learned templates selected by the Potts states:
-$\mu(h)=b+\sum_{j=1}^{m}W_{j}^{(h_{j})}$ (4)
-This creates a structured and factored representation in visible space.
-Each hidden unit encodes one of $q$ possible modes, and the overall visible configuration is a compositional sum of these latent components.
-This modularity enables reuse of latent features across samples, disentangled latent representation, and combinatorial expressivity across $q^{m}$ configurations
-The set of all compositional means $\mu(h)$ forms a structured lattice in $\mathbb{R}^{n}$. The resulting energy landscape has the following properties:
-  • Locally linear: Each mode defines a Gaussian with fixed covariance
-  • Globally discrete: The mode structure is defined by a combinatorial Potts lattice
-  • Modular: Latent units contribute independently to the visible pattern
-This structure allows GM-RBMs to model multimodal distributions over continuous space with precise symbolic control via latent discrete selections.
+Restricted Boltzmann Machines (RBMs) are energy-based models with an undirected bipartite graph structure with visible and hidden units and a subset of the broader fully connected Boltzmann machines.
+The lack of intralayer connections enables training tractability through parallel updates of all units within each layer (Block Gibbs updates) through a biased stochastic gradient estimator for log-likelihood (Contrastive Divergence) (Hinton, 2002; Hinton and Salakhutdinov, 2006).
+Although powerful, the strictly binary units in hidden and visible units make them non-ideal for dealing with multivalued data.
+One approach to extend their use to continuous data is to replace the visible units with Gaussian units, termed Gaussian–Bernoulli RBMs (GB-RBMs) (Liao et al., 2022).
+However, their binary hidden units often struggle with inherently categorical, mutually exclusive factors.
+We address this mismatch by replacing each binary hidden unit with a one-of-$q$ categorical unit (Potts), yielding the Gaussian–Multinoulli RBM (GM-RBM): Gaussian visibles paired with $q$-state latent hidden units.
+Conceptually, this keeps the RBM’s simplicity but aligns inductive bias with categorical structure, effectively allowing the model to capture inherent, underlying categorical relationships.
+Prior work tended to retain binary latents as they fit existing tooling and samplers; in contrast, categorical slots raise practical issues (e.g., intra-slot degeneracy) and make fair comparisons challenging.
+In order to separate architectural effects from raw capacity, we design comparison protocols and clarify the challenges associated with them.
+Our contributions are threefold: (i) a drop-in Potts hidden layer that preserves tractable conditionals while retaining the standard RBM training pipeline; (ii) comparison protocols (capacity-matched and parameter-matched) that isolate the effect of categorical slots; and (iii) empirical results showing that, with equal negative-phase budgets and pure block Gibbs updates, increasing $q$ consistently improves image quality (FID) and hetero-associative recall, closing much of the gap to GB-RBMs that require costlier Gibbs–Langevin steps (Liao et al., 2022).
+The key message being a minimal architectural change yields disproportionate gains due to the q-slot augmented latent: sharper posteriors, more interpretable codes, and stronger retrieval and generation while utilizing similar compute resources.
+2 Background and Motivation
+2.1 Background
+Boltzmann machines (BM) are a class of energy-based models utilizing binary units ($\{-1,1\}$) that sample from a Boltzmann probability distribution over $x=(v,h)$ given by:
+$p_{\theta}(x)=\frac{1}{Z(\theta)}\exp\!\big(-E_{\theta}(x)\big),\qquad Z(\theta)=\sum_{x}\exp\!\big(-E_{\theta}(x)\big).$
+where $\theta$ is a set of biases ($a_{i}$) and weights ($W_{ij}$). (Smolensky, 1986; Hinton, 2002) The energy function for a BM is given by:
+$E_{\theta}(x)=-\sum_{i}a_{i}x_{i}-\tfrac{1}{2}\sum_{i\neq j}W_{ij}x_{i}x_{j},\qquad W_{ij}=W_{ji},\;W_{ii}=0.$
+Log-likelihood learning decomposes into a positive (data) and negative (model) phase:
+$\frac{\partial}{\partial\theta}\log p_{\theta}(v)=\mathbb{E}_{p(h\mid v)}\!\Big[-\tfrac{\partial E(v,h)}{\partial\theta}\Big]-\mathbb{E}_{p(v\mid h)}\!\Big[-\tfrac{\partial E(v,h)}{\partial\theta}\Big],$
+which yields the classic correlation-matching rules:
+$\frac{\partial}{\partial W_{ij}}\log p_{\theta}(v)=\langle x_{i}x_{j}\rangle_{\text{data}}-\langle x_{i}x_{j}\rangle_{\text{model}},\qquad\frac{\partial}{\partial a_{i}}\log p_{\theta}(v)=\langle x_{i}\rangle_{\text{data}}-\langle x_{i}\rangle_{\text{model}}.$
+However, correlation-matching rules are intractable for most real-world data (Liao et al., 2022).
+One way to overcome this is to use the Restricted Boltzmann machine: Restricted Boltzmann machines split the graph into a bipartite structure with visible units $v$ and hidden units $h$, removing within-layer edges so that the joint probability distribution can be broken down into simpler conditional probabilities enabling block Gibbs sampling.
+For binary visibles $v\in\{0,1\}^{n}$ and binary hiddens $h\in\{0,1\}^{m}$,
+$E(v,h)=-b^{\top}v-c^{\top}h-v^{\top}Wh,$
+$p(h_{j}=\textbf{1}\mid v)=\sigma\big((W^{\top}v)_{j}+c_{j}\big),\qquad p(v_{i}=\textbf{1}\mid h)=\sigma\big((Wh)_{i}+b_{i}\big),$
+and the learning updates specialize to moment matching between the two layers,
+$\Delta W\propto\mathbb{E}[vh^{\top}]_{\text{data}}-\mathbb{E}[vh^{\top}]_{\text{model}},\quad\Delta b\propto\mathbb{E}[v]_{\text{data}}-\mathbb{E}[v]_{\text{model}},\quad\Delta c\propto\mathbb{E}[h]_{\text{data}}-\mathbb{E}[h]_{\text{model}}.$
+Contrastive divergence(CD) approximates model expectations by alternating $h\!\sim\!p(h\mid v)$ and $v\!\sim\!p(v\mid h)$ for a $k$ steps of data CD-$k$ or for a persistence chain (persistent CD).
+Usually, both hidden and visible units are taken as binary, but for continuous data it is natural to use Gaussian visible units (Liao et al., 2022).
+In the Gaussian–Bernoulli RBM (Hinton), visibles are Gaussian with diagonal variance $\sigma^{2}$ and hiddens remain binary.
+A convenient parameterization is:
+$E(v,h)=\sum_{i}\frac{(v_{i}-\mu_{i})^{2}}{2\sigma_{i}^{2}}-\sum_{i,j}\frac{v_{i}}{\sigma_{i}^{2}}W_{ij}h_{j}-\sum_{j}b_{j}h_{j},$
+which gives closed-form conditionals used inside CD,
+$p(v\mid h)=\mathcal{N}\!\big(\mu+Wh,\ \mathrm{diag}(\sigma^{2})\big),\qquad p(h_{j}=1\mid v)=\text{Sigmoid}\!\Big(\big[W^{\top}(v\oslash\sigma^{2})\big]_{j}+b_{j}\Big).$
+The corresponding stochastic-gradient updates keep the same positive/negative structure but respect the visible scaling,
+$\Delta\bm{W}\propto\Big\langle\left(\frac{\mathbf{v}}{\bm{\sigma}^{2}}\right)\,\mathbf{h}^{\top}\Big\rangle_{\text{data}}-\Big\langle\left(\frac{\mathbf{v}}{\bm{\sigma}^{2}}\right)\,\mathbf{h}^{\top}\Big\rangle_{\text{model}},$
+$\Delta\bm{\mu}\propto\Big\langle\frac{\mathbf{v}-\bm{\mu}}{\bm{\sigma}^{2}}\Big\rangle_{\text{data}}-\Big\langle\frac{\mathbf{v}-\bm{\mu}}{\bm{\sigma}^{2}}\Big\rangle_{\text{model}}$
+$\Delta\bm{b}\propto\langle\bm{h}\rangle_{\text{data}}-\langle\bm{h}\rangle_{\text{model}}\qquad$
+In this work we extend the role of the binary hidden variables by replacing each Bernoulli with a one-of-$q$ categorical (Potts) slot, aligning the latent prior with mutually exclusive factors while preserving the RBM’s block-sampling tractability and the Gaussian visible layer.
+2.2 Motivating Principles
+Many perceptual and symbolic factors are naturally categorical and mutually exclusive.
+Approximating such structure with many Bernoulli latents (as in a GB–RBM) forces variants to be represented by co–activating subsets of units, which encodes information across the hidden layer and yields ambiguous codes.
+The Gaussian–Multinoulli RBM (GM–RBM) encodes each factor as a one-of-$q$ slot.
+With $m$ slots and hidden configuration $h=(h_{1},\dots,h_{m})$ where $h_{j}\in\{1,\dots,q\}$, the visible conditional remains Gaussian,
+$p(v\mid h)=\mathcal{N}\!\Big(b+\sum_{j=1}^{m}W^{(h_{j})}_{:,j},\,I\Big),$
+so each chosen state contributes a single template vector and the mean is a sum of selected templates.
+This preserves the RBM’s locally linear structure encoding the continuous variables using the set of latent discrete variables.
+A key advantage of the Potts node is that each slot has its own weight vector, so swapping slots does not alter the previous slot’s learned weight.
+In practice, this yields sharper posteriors and more degrees of freedom than simply enforcing a one-hot encoding across multiple Bernoullis, which intrinsically lack intra-layer coupling.
+GM–RBM is a modification where we replace binary hidden units with categorical slots but retain Gaussian visibles and standard RBM conditionals.
+We assess GM–RBM on hetero-associative recall and image modeling.
+To highlight the intrinsic strength of the GMRBM to the GBRBM, we develop comparison protocols that are capacity-matched (hidden nodes) and parameter-matched (weights), which appear later.
+3 Theoretical Foundations of the Gaussian–Multinoulli RBM (GM-RBM)
+The Gaussian–Multinoulli RBM (GM-RBM) replaces binary hidden units with discrete $q$-state categorical variables while keeping a continuous visible layer.
+This yields a combinatorial latent space with simple, closed-form conditionals.
+3.1 Notation
+Let $v\in\mathbb{R}^{n}$ be the visible vector.
+The hidden code is $h=(h_{1},\dots,h_{m})$ with $h_{j}\in\{1,\dots,q\}$.
+Parameters are: visible bias $b\in\mathbb{R}^{n}$; hidden bias $c_{j,k}\in\mathbb{R}^{m}$; and state-specific templates $W^{(k)}_{:,j}\in\mathbb{R}^{n}$.
+Define the conditional mean
+$\mu(h)=b+\sum_{j=1}^{m}W^{(h_{j})}_{:,j}$
+3.2 Energy, Joint, and Conditionals
+The energy is
+$E(v,h)=\tfrac{1}{2}\sum_{i=1}^{n}(v_{i}-b_{i})^{2}-\sum_{j=1}^{m}c_{j,h_{j}}-\sum_{i=1}^{n}\sum_{j=1}^{m}W^{(h_{j})}_{ij}\,v_{i}.$
+Completing the square gives $E(v,h)=\tfrac{1}{2}\lVert v-\mu(h)\rVert_{2}^{2}+K(h)$ with $K(h)=\tfrac{1}{2}(\lVert b\rVert_{2}^{2}-\lVert\mu(h)\rVert_{2}^{2})-\sum_{j}c_{j,h_{j}}$.
+The joint is $p(v,h)\propto\exp(-E(v,h))$.
+The conditionals are
+$p(v\mid h)=\mathcal{N}\big(\mu(h),\bm{1}\big)$
+$\qquad p(h_{j}=k\mid v)=\frac{\exp\big(c_{j,k}+(W^{(k)}_{:,j})^{\top}v\big)}{\sum_{k^{\prime}=1}^{q}\exp\big(c_{j,k^{\prime}}+(W^{(k^{\prime})}_{:,j})^{\top}v\big)}=\text{Softmax}{(c_{j,k^{\prime}}+(W^{(k^{\prime})}_{:,j})^{\top}v\big)}$
+3.3 Architecture and Special Cases
+Each slot contributes one of $q$ templates, so the codebook $\{\mu(h):h\in\{1,\dots,q\}^{m}\}$ has size $q^{m}$.
+When $q=2$ and parameters are tied as $W^{(1)}_{:,j}-W^{(2)}_{:,j}=\widetilde{W}_{:,j}$ and $c_{j,1}-c_{j,2}=\widetilde{b}_{j}$ with a corresponding recentering of $b$, the GM-RBM reduces to a Gaussian–Bernoulli RBM.
+There is no requirement that $q$ be even; experiments may vary $q$ according to task.
+It is also possible to use usual variance reweighting as opposed to unit variance used in the GB-RBM.
+3.4 Parameter Count and Fair Comparison Protocols
+GM-RBM has $n$ parameters for $b$, plus $mq$ for $c$, plus nmqnmq for $W$, for a total of n+mq(1+n)n+mq\,(1+n).
+A GB-RBM with $m^{\prime}$ binaries has $n+m^{\prime}+nm^{\prime}$.
+We evaluate under two protocols:
+  • Parameter-matched: Match the number of latent assignments by setting $m^{\prime}\approx m\,\log_{2}q$, since $|\mathcal{H}_{\text{GM}}|=q^{m}$ and $|\mathcal{H}_{\text{GB}}|=2^{m^{\prime}}$.
+Essentially, choose $m^{\prime}$ so total parameters are comparable across models
+  • Capacity-matched: Choose $|\mathcal{H}_{\text{GB}}|$ in such a way that it matches the total number of learned representation for a given $|\mathcal{H}_{\text{GM}}|$
+These regimes separate architectural effects (slot exclusivity and templating) from raw codebook size.
+3.5 Learning and Negative-phase Sampling
+The log-likelihood gradient satisfies
+$\tfrac{\partial}{\partial\theta}\log p(v)=\mathbb{E}_{p(h\mid v)}\Big[-\tfrac{\partial E(v,h)}{\partial\theta}\Big]-\mathbb{E}_{p(v\mid h)}\Big[-\tfrac{\partial E(v,h)}{\partial\theta}\Big],\quad\theta\in\{b,c,W\}.$
+We approximate the model expectation with short Markov chains.
+Block Gibbs.
+Alternate $h\sim p(h\mid v)$ using the per-slot softmax and $v\sim p(v\mid h)=\mathcal{N}(\mu(h),\mathbf{1})$. The visible draw is exact and parameter free.
+Gibbs with visible Langevin.
+Some implementations replace the exact Gaussian draw with an unadjusted Langevin step using $\nabla_{v}\log p(v\mid h)=\mu(h)-v$ (Liao et al., 2022):
+$v_{t+1}=v_{t}+\tfrac{\varepsilon^{2}}{2}\big(\mu(h_{t})-v_{t}\big)+\varepsilon\,\xi_{t},\quad\xi_{t}\sim\mathcal{N}(0,\mathbf{1}),\ \varepsilon>0.$
+This introduces a stepsize–dependent discretization error; for fixed $h_{t}$, multiple Langevin steps only approximate the exact conditional.
+In many Gaussian RBM variants, each visible dimension is allowed its own variance, so that $p(v\mid h)=\mathcal{N}(\mu(h),\Sigma)$ with a diagonal covariance $\Sigma=\operatorname{diag}(\sigma^{2})$.
+In this case, the gradient becomes $\Sigma^{-1}(\mu(h_{t})-v_{t})$, and the drift term is modified to $(\mu(h_{t})-v_{t})/\sigma^{2}$ (applied elementwise).
+Intuitively, inserting a visible Langevin move between hidden-layer updates allows hidden units to exchange information indirectly through the visible layer, encouraging more cooperative encodings and improving the model’s ability to capture complex dependencies.
+In contrast to the expensive Langevin update, we use exact block Gibbs updates for the GM-RBM.
+The reason we utilize this less expensive block Gibbs update is that the visible Langevin step is merely an approximate sampler for the same conditional distribution $p(v\mid h)$.
+It does not inject additional information into the model, but instead introduces stepsize-dependent bias at an extra computational cost.
+Meanwhile, the multi-state multinoulli latent variables already enable this information to be represented and shared across hidden units through standard block Gibbs updates, making an additional visible Langevin refinement unnecessary in our setting.
+To ensure fair comparison to GB-RBM baselines, we explicitly report whether a visible Langevin step is used, the stepsize, the number of steps, and whether chains are persistent.
+Sampler cost, mixing, and our choice.
+The visible Langevin variant adds at least one extra update and a stepsize hyperparameter per negative step, which increases compute relative to an exact Gaussian draw.
+In our setting the Gibbs visible update samples exactly from $N(\mu(h),\mathbf{1})$ with a single noise vector and no stepsize tuning.
+The GM-RBM did was able to achieve quality results without resorting to the more expensive Langevin update.
+This choice also highlights a modeling claim of this work: with slot-wise categorical latents, GM-RBM exhibits fast mixing under basic Gibbs updates without relying on heavier samplers.
+We report chains and effective sample diagnostics to substantiate this point.
+3.6 Key Properties
+  • Locally linear: given $h$, $p(v\mid h)$ is Gaussian with fixed covariance.
+  • Globally discrete: the means form a finite codebook indexed by discrete slots.
+  • Modular: slots contribute additively in $\mu(h)$ and independently in $p(h\mid v)$.
 4 Hetero-associative Memory
-Hetero-associative memory refers to a system’s capability to learn paired associations between distinct patterns, allowing the retrieval of a target pattern (e.g., a response word) when presented with a corresponding stimulus (e.g., a cue word) [11, 19].
-This concept, rooted in cognitive modeling and neural computation, was further explored by Hinton in 1981 [2] and later employed in language-related tasks using Gaussian-Bernoulli Restricted Boltzmann Machines (GB-RBMs) [27].
+Hetero-associative memory refers to a system’s capability to learn paired associations between distinct patterns, allowing the retrieval of a target pattern (e.g., a response word) when presented with a corresponding stimulus (e.g., a cue word) (Kosko, 1988; Morales and Pineda, 2024).
+This concept, rooted in cognitive modeling and neural computation, was further explored by Hinton in 1981 Anderson and Hinton (1981) and later employed in language-related tasks using Gaussian-Bernoulli Restricted Boltzmann Machines (GB-RBMs) (Tsutsui and Hagiwara, 2019).
 However, the binary hidden units of GB-RBMs impose a hard ceiling on representational capacity.
@@
 4.1 Experimental Setup
-4.1.1 Dataset
-We replicated the experimental setup in Tsutsui and Hagiwara, constructing a word-pair dataset representing conceptual relationships (e.g., "apple is-a fruit") [27].
-We randomly selected pairs from WordNet [28], excluding compound or incomplete entries and sampled 500-3,000 word pairs to create smaller datasets for training and testing scalability.
-Each word pair is treated as a directional association task and encoded as a concatenated embedding vector [17], mirroring the hetero-associative memory objective of recalling a target concept from a given stimulus.
-4.1.2 Preprocessing
-We trained a 200-dimensional Continuous Bag of Words (CBOW) Word2Vec model [18] on the word-pair dataset (100 iterations, window size 5, no frequency cutoff).
+We replicated the experimental setup in Tsutsui and Hagiwara, constructing a word-pair dataset representing conceptual relationships (e.g., "apple is-a fruit") (Tsutsui and Hagiwara, 2019).
+We randomly selected pairs from WordNet (Princeton University (2010)), excluding compound or incomplete entries and sampled 500-3,000 word pairs to create smaller datasets for training and testing scalability.
+Each word pair is treated as a directional association task and encoded as a concatenated embedding vector (Mikolov et al., 2013), mirroring the hetero-associative memory objective of recalling a target concept from a given stimulus.
+We trained a 200-dimensional Continuous Bag of Words (CBOW) Word2Vec model (Mikolov et al., 2013) on the word-pair dataset (100 iterations, window size 5, no frequency cutoff).
 We normalized each vector dimension to zero mean and unit variance to mitigate small-magnitude issues.
 For each stimulus-response pair, embeddings were concatenated to form a 400-dimensional input vector for the RBM’s visible layer.
-4.1.3 Computational Setup
 We used a compute node (Intel Xeon Gold 6154 ×2, 512 GB RAM, NVIDIA Tesla P40) running Red Hat Enterprise Linux 8.
@@
 Experiments were automated via a modular framework for data loading, configuration, visualization, and checkpointing.
-4.1.4 Training
+4.2 Training
 Our GM-RBM learned to associate stimulus–response word embeddings.
 Datasets comprised semantically related word pairs (e.g., doctor–nurse, sun–light).
-Following a similar procedure to Tsutsui and Hagiwara [27], we trained a CBOW Word2Vec model [23] on these two-word sentences to capture in-domain semantics, producing 200-dimensional embeddings (100 iterations, window size 5, no frequency cutoff).
+Following a similar procedure to Tsutsui and Hagiwara Tsutsui and Hagiwara (2019), we trained a CBOW Word2Vec model (Řehůřek and Sojka, 2010) on these two-word sentences to capture in-domain semantics, producing 200-dimensional embeddings (100 iterations, window size 5, no frequency cutoff).
 We normalized each embedding dimension to zero mean and unit variance to avoid numerical instability.
@@
 The second was a similar sweep where $q$ was held constant while the number of hidden units themselves was increased.
-Both sweeps were done across varying dataset sizes, measuring the accuracy of the recall in pairs embedded in Word2Vec [18].
+Both sweeps were done across varying dataset sizes, measuring the accuracy of the recall in pairs embedded in Word2Vec (Mikolov et al., 2013).
 Each marker in Figures 1 and 2 denotes an independently trained model using identical early-stopping criteria.
@@
 The original GB-RBM outperforms the GM-RBM for $q=2$, possibly due to the lack of this constraint.
-Despite the GM-RBMs being trained using the simpler Gibbs update, the $q=4,6,8,10$ GM-RBM models all vastly outperform the GB-RBM, which used the more expensive Gibbs-Langevin update [15].
+Despite the GM-RBMs being trained using the simpler Gibbs update, the $q=4,6,8,10$ GM-RBM models all vastly outperform the GB-RBM, which used the more expensive Gibbs-Langevin update (Liao et al., 2022).
 Since constraining the total number of parameters to be identical might obfuscate the potential of both models, we swept the number of hidden nodes and dataset sizes as shown in Figure 2.
@@
 Notably, $q=10$ consistently outperforms lower-$q$ configurations even for large $N$.
-To keep the model capacity fixed across different Potts-state configurations, we solve for the number of hidden units $n_{h}$ by dividing the total budget of weight parameters $n_{w}$ by the size of each hidden Potts spin $q$.
+To keep the model capacity fixed across different Potts‐state configurations, we solve for the number of hidden units $n_{h}$ by dividing the total budget of weight parameters $n_{w}$ by the size of each hidden Potts spin $q$.
 Since each hidden unit with $q$ Potts states contributes $q$ weight vectors of length equal to the number of visible units $n_{v}$, then number of hidden units is given by
 $n_{h}\texttt{=}\lceil\frac{n_{w}}{n_{v}}\times q\ \rceil$ (1)
-We chose $\mathrm{n_{w}}=800{,}000$, ensuring that the binary case $q=2$ matched the optimal hidden-unit count of 1,000 reported by Tsutsui and Hagiwara [27]for the GB-RBM.
+We chose $\mathrm{n_{w}}=800{,}000$, ensuring that the binary case $q=2$ matched the optimal hidden‐unit count of 1,000 reported by Tsutsui and Hagiwara Tsutsui and Hagiwara (2019)for the GB-RBM.
 This ensures that, as we increase the number of Potts states $k$, we reduce the number of hidden units proportionally so that every model has the same total number of parameters.
@@
 4.3.2 Hidden nodes sweep
-Figure 2: Retrieval accuracy on a semantic hetero-associative memory task as a function of hidden-layer size for different model variants and dataset sizes.
+Figure 2: Retrieval accuracy on a semantic hetero‐associative memory task as a function of hidden‐layer size for different model variants and dataset sizes.
 (a) GB-RBM (Gibbs–Langevin Update), (b) GM-RBM with $q=2$ (Gibbs Update), and (c) GM-RBM with $q=4$ (Gibbs Update).
-Each curve corresponds to a different number of training word-pair examples ($N$).
+Each curve corresponds to a different number of training word‐pair examples (NN).
 In order to showcase how varying the hidden-layer dimensionality affects retrieval performance across dataset sizes, we swept the number of hidden units from 500 to 3000 while fixing the state cardinality $q$.
@@
 5 Auto-associative memory
-As a "proof-of-concept" to demonstrate the generative capability of our GM-RBM, we performed a replication of the generative experiments from the original GB-RBM paper [15].
+As a "proof-of-concept" to demonstrate the generative capability of our GM-RBM, we performed a replication of the generative experiments from the original GB-RBM paper Liao et al.
+(2022).
 On MNIST, we trained for 500 epochs, and on CelebA for 100 epochs.
@@
 5.1 Experimental setup
-5.1.1 Datasets
-We target two key datasets to showcase the GM-RBM’s generative abilities: MNIST (28×28 grayscale handwritten digits) [13] and CelebA (Center-cropped and resized RGB facial images at 64×64 resolution)  [16].
-5.1.2 Preprocessing
-All image datasets are normalized to zero mean and unit variance per channel [15].
-CelebA images are center-cropped to 140×140 before downsampling to 64×64 [16].
-The GMM datasets are constructed using predefined priors and covariance structures, allowing for exact density visualization and sampling [15].
+We target two key datasets to showcase the GM-RBM’s generative abilities: MNIST (28×28 grayscale handwritten digits) (LeCun et al., 1998) and CelebA (Center-cropped and resized RGB facial images at 64×64 resolution)  (Liu et al., 2015).
+All image datasets are normalized to zero mean and unit variance per channel (Liao et al., 2022).
+CelebA images are center-cropped to 140×140 before downsampling to 64×64 (Liu et al., 2015).
+The GMM datasets are constructed using predefined priors and covariance structures, allowing for exact density visualization and sampling (Liao et al., 2022).
 For each dataset, we generate training and evaluation splits using standard protocols.
-5.1.3 Computational resources
 Experiments were performed on a dedicated compute node equipped with dual Intel Xeon Gold 5218 CPUs, 512 GB of RAM, and eight NVIDIA RTX 6000 GPUs connected via NVLink.
@@
 (a) Selected samples from MNIST
-(b) Generation of CelebA images from random noise
-Figure 3: Sampled results from GM-RBM of MNIST and CelebA datatsets
-We observe that the $q=4$ GM-RBM begins to generate visually identifiable face/digit samples with an order of magnitude lower number of epochs compared to the GB-RBM trained with Gibbs–Langevin sampling (See Table 2.
-Although a fully controlled head-to-head comparison under identical hyperparameter budgets (see Table 2) requires further investigation (See Section 6.1, it is important to emphasize two key distinctions.
-Our GM-RBM relies solely on conventional Gibbs sampling, whereas Hinton’s GB-RBM employs the more computationally expensive Gibbs–Langevin procedure [15].
-This ability is most likely due to what the rapid mixing of Potts models compared to binary Boltzmann machines [22, 31, 32].
+(b) Generation of CelebA images from random noise with $q=4$
+Figure 3: Sampled results from GM-RBM of MNIST and CelebA datasets.
+We observe that the $q=4$ GM-RBM begins to generate visually identifiable face/digit samples (See Figure 3(a) and 3(b)) with an order of magnitude lower number of epochs compared to the GB-RBM trained with Gibbs–Langevin sampling (See Table 2.
+Although a fully controlled head-to-head comparison under identical hyperparameter budgets (see Table 2) requires further investigation (See Section 6.1), it is important to emphasize two key distinctions.
+Our GM-RBM relies solely on conventional Gibbs sampling, whereas Hinton’s GB-RBM employs the more computationally expensive Gibbs–Langevin procedure (Liao et al., 2022).
+This ability is most likely due to the rapid mixing of Potts models compared to binary Boltzmann machines (Potts, 1952; Whitehead et al., 2023; Wu, 1982).
 In addition, we maintain the same total number of parameters while increasing the number of Potts states.
-This reduction in both sampling complexity and overall training budget underscores the richer latent representational capacity afforded by Potts-style hidden units, which can capture more nuanced multimodal structure with fewer resources.
-The stark difference in hyperparameters was chosen specifically to highlight the GM-RBM’s ability to achieve high-quality samples under a dramatically reduced resource budget, showcasing its practical efficiency.
 Table 2: Hyperparameter settings for GB-RBM and GM-RBM on MNIST and CelebA
-Hyperparameter MNIST CelebA
-GB-RBM  [15] GM-RBM GB-RBM [15] GM-RBM
+MNIST CelebA
+Hyperparameter GB-RBM GM-RBM GB-RBM GM-RBM
 Number of states 2 4 2 4
@@
 Epochs trained 3000 500 10000 100
+This reduction in both sampling complexity and overall training budget underscores the richer latent representational capacity afforded by Potts-style hidden units, which can capture more nuanced multimodal structure with fewer resources.
+The stark difference in hyperparameters was chosen specifically to highlight the GM-RBM’s ability to achieve high-quality samples under a dramatically reduced resource budget, showcasing its practical efficiency.
+5.2.1 Quantitative sample quality (FID)
+Table 3: FID (↓) under parameter-matched budgets. Best overall in bold.
+Model Potts States FID
+GM–RBM $q=2$ 67.08
+GM–RBM $q=4$ 56.09
+GM–RBM $q=6$ 53.07
+GB–RBM N/A 60.06
+To complement the qualitative samples, we report Fréchet Inception Distance (Heusel et al., 2018) (FID; ↓ is better) under a capacity-matched protocol ($m^{\prime}\approx m\log_{2}q$), holding negative-phase budgets fixed across models.
+Results show that increasing $q$ consistently improves GM–RBM sample quality under a pure Gibbs sampler, narrowing the gap to GB–RBM with Gibbs–Langevin despite the latter’s higher per-step compute.
+With the Gibbs-only updates, GM–RBM (q=6q{=}6) outperforms GB–RBM by $\approx$7 FID points (53.07 vs 60.06).
+With a more expensive update, it is possible to lower the FID score.
+It is important to note that the CelebA dataset does not necessarily have clear relational qualities, meaning that these scores are really showcasing the mixing ability of the different sampling algorithms and models.
 6 Limitations and future work
-While the Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM) demonstrates improved expressiveness and recall performance over the binary GB-RBM, key limitations and future work remain.
+While the Gaussian–Multinoulli RBM (GM–RBM) demonstrates stronger discrete expressiveness and robust recall under fair matching, important limitations and next steps remain.
 6.1 Limitations of the current model
 1.
-Sampling constraints: In the current implementation, only Gibbs sampling is supported for the Potts-based GM-RBM.
-While this allows for tractable inference over discrete states, the absence of Langevin or hybrid Gibbs-Langevin samplers limits potential gains in sample quality and mixing speed.
-As $q$ increases, the discrete state space grows, and relying solely on categorical sampling may lead to slower convergence or poor exploration of the energy landscape in more complex settings.
+Sampling constraints and rationale.
+Our GM–RBM uses pure block Gibbs (exact Gaussian visible draw + per-slot softmax posteriors).
+We intentionally avoid visible-space Langevin during training because it adds step-size hyperparameters, extra updates, and discretization error, increasing compute relative to a single exact Gaussian draw.
 2.
-Evaluation domain: The current evaluation focuses on word pair associations using Word2Vec embeddings.
-The generalization of GM-RBMs to other domains, such as vision, audio, or time series data, remains an open question and will require further validation.
-These limitations present valuable directions for future research, including scaling GM-RBMs to deeper architectures, optimizing discrete sampling methods, and exploring broader applications beyond associative memory tasks.
+Evaluation scope and scaling.
+Most of our positive results are on hetero-associative recall with in-domain Word2Vec embeddings and proof-of-concept image generation.
+This leaves open: (a) robustness to stronger or off-the-shelf embeddings (e.g., large CBOW/skip-gram trained on broader corpora), (b) transfer to other modalities (audio, trajectories, time series), and (c) deeper stacks (e.g., GM front-ends for DBMs/DBNs).
+We also observe retrieval degradation as $N$ grows in some settings; diagnosing whether this arises from chain mixing, capacity-matching choices, or dataset effects will require ablations (burn-in length, persistence, $q$ vs.
+$m$, negative-phase steps).
 3.
-Epochs and training: Currently, the GM-RBM is trained for a limited number of epochs (500 for MNIST, 100 for CelebA) to achieve initial convergence.
-While this rapid training demonstrates efficiency, it may not capture long-term stability or the potential benefits of extended training schedules.
-Optimal epoch budgets, early-stopping criteria, and the impact of prolonged training on sample diversity and model robustness are areas for further exploration.
+Training budget and stability.
+Our generative experiments used short schedules (e.g., 500 epochs on MNIST, 100 on CelebA) to highlight practicality under a reduced budget.
+Longer training with early-stopping and variance monitoring could reveal stability regimes, impacts on sample diversity, and when heavier negative-phase samplers pay off.
 6.2 Broader impact
 6.2.1 Potts units in energy transformers
-Energy Transformers (ETs) minimize a global energy via recurrent updates and have shown success in tasks such as image inpainting [9].
-Replacing their binary hidden units with $q$-state Potts variables raises the latent capacity from $2^{H}$ to $q^{H}$, yielding:
-$E_{\mathrm{mem}}(h)=-\sum_{\mu=1}^{M}\sum_{i=1}^{H}\delta(h_{i},\xi_{i}^{\mu})$ (6)
-where ${\xi^{\mu}}$ are stored patterns.
-Larger $q$ reduces overlap between attractors, improving convergence selectivity.
-Embedding Gaussian–Potts RBMs into ETs may boost robustness and facilitate neuromorphic implementations; an empirical study of these hybrids is ongoing.
-6.2.2 Extending DBMs to continuous data
-Deep Boltzmann Machines (DBMs) stack Bernoulli–Bernoulli RBMs to learn hierarchical features via layer-wise Gibbs sampling and contrastive divergence [3, 24].
-Their reliance on binary visibles limits real-valued inputs.
-As a front end, we propose a Gaussian–Potts energy:
-$E(v,h)=\sum_{i=1}^{D}\frac{(v_{i}-a_{i})^{2}}{2\sigma_{i}^{2}}-\sum_{i,j}W_{ij}(v_{i},h_{j})-\sum_{j}b_{h_{j}},$ (7)
-with $h\_j\in{1,\dots,q}$. This encoder maps $\mathbb{R}^{D}$ into a $q^{H}$-state latent space, preserving DBM training while improving stability and expressiveness.
-6.3 Bernoulli-Variable models and Potts benefits
-Many foundational generative architectures rely on binary or Bernoulli units, including classical models like Bernoulli RBMs [6], Gaussian–Bernoulli RBMs [15], Deep Belief Networks [24], Deep Boltzmann Machines [3, 24], and Hopfield Networks.
-More recent advancements leveraging binary sampling mechanisms include Variational Autoencoders (VAEs) [5] with Bernoulli latents, Gumbel–Softmax VAEs [21] that provide differentiable relaxations for discrete sampling, Autoregressive Models such as PixelCNN [30] and PixelRNN [29] for sequential binary pixel generation, Binarized Neural Networks (BNNs) that constrain weights and activations to binary values, Latent Bernoulli Autoencoders [5] designed explicitly for compact binary representations, Spiking Neural Networks (SNNs) employing discrete spike events for communication, and Discrete Denoising Diffusion Models (D3PMs) [8] that extend diffusion processes to discrete data through multinomial transition kernels.
-Upgrading these binary units to Potts variables further increases representational granularity and sharpens attractor basins, significantly reducing interference among stored patterns and enabling richer discrete structures, ultimately enhancing the capability and expressiveness of associative and generative models.
-6.4 Hardware-Efficient Implementation
-Many of these models can be deployed with minimal overhead by leveraging sparse or event-driven representations (e.g., SPAD arrays [31]) and one-hot encoding of discrete states.
-Binary and Potts variables map naturally to bitwise logic and lookup tables, enabling FPGA, ASIC, and neuromorphic implementations with low memory footprint and energy consumption.
+Energy Transformers minimize a global energy via recurrent updates.
+Replacing binary hidden units with $q$-state Potts slots increases latent capacity from $2^{H}$ to $q^{H}$ and reduces attractor overlap:
+$E_{\mathrm{mem}}(h)\;=\;-\sum_{\mu=1}^{M}\sum_{i=1}^{H}\delta(h_{i},\xi_{i}^{\mu}).$ (2)
+We hypothesize improved selectivity/robustness under higher $q$ and are empirically exploring GM–RBM components as ET memories; careful comparisons will match either capacity or parameter budgets as above.
+(Schröder et al., 2024; Hoover et al., 2023)
+As a front-end to DBMs, a Gaussian–Potts encoder maps $\mathbb{R}^{D}$ into a $q^{H}$ latent space while preserving layer-wise Gibbs training.
+Future work should examine stability under deeper stacks and quantify benefits over increased Bernoulli width at matched capacity.
+(Morales and Pineda, 2024; Oh et al., 2022)
+Many foundational generative families, RBMs, DBNs/DBMs, Hopfield-style memories, VAEs with Bernoulli latents, Gumbel-Softmax relaxations, autoregressive binarized pixel models, BNNs, SNNs, and discrete diffusion—lean on binary sampling.
+Swapping binary units for Potts slots increases representational granularity and sharpens attractor basins, which can reduce interference among stored patterns at equal capacity.
+A broad, capacity- or parameter-matched survey across these families is an open, high-impact direction (Kosko, 1988).
+Binary and Potts one-hot codes map naturally to LUTs and bitwise logic.
+Sparse/event-driven readout and lightweight on-chip softmax enable efficient FPGA/ASIC/neuromorphic realizations; exploring SPAD-assisted annealing or mixed-signal implementations for categorical slots is promising future work (Whitehead et al., 2023).